OpenStudy (anonymous):

Statistics: Let X have the density function f(x) = cx(1-x) for 0<=x<=1 and f(x)=0 otherwise. Find c. Do I integrate the function between the values 0 and 1?

OpenStudy (turingtest):

why not?\[\int_{0}^{1}cx(1-x)dx\]

OpenStudy (turingtest):

...=1 I guess, right?

OpenStudy (anonymous):

How did you get that? I brought the c outside what's being integrated and then multiplied out the function. So now I have \[c \int\limits_{0}^{1}(x^2/2)-(x^3/3)dx\] Then subbing in 1 and 0 I ended up with c*1/6. Have I done that much correctly?

OpenStudy (turingtest):

you have already integrated it, but you left the integral sign o-0 ?

OpenStudy (anonymous):

Sorry. That's a typo :D

OpenStudy (turingtest):

\[\int_{0}^{1}cx(1-x)dx=1\]\[c\int x-x^2dx=1\]\[c(\frac12x^2-\frac13x^3)|_{0}^{1}=1\]

OpenStudy (anonymous):

right you should get c(1/2 - 1/3) =1, then solve for c

OpenStudy (turingtest):

\[c(\frac12-\frac13)=c\frac16=1\implies c=6\]

OpenStudy (anonymous):

Oh right, sorry Turing, I thought you were getting 1 as the answer for the integration. :D

OpenStudy (anonymous):

So that's a rule that applies to all functions in pdf, is it? You just let them equal to 1?

OpenStudy (turingtest):

for a probability density functions \[\int_{-\infty}^{\infty}f(x)dx=1\]since the function is zero everywhere except the interval [0,1], we only needed to integrate along those bounds

OpenStudy (turingtest):

if you have never seen that formula before, try to think about the logic behind it.

OpenStudy (anonymous):

I'm guessing it's equal to one because of the rule in statistics that all probabilities must be between 0 and 1?